OpenStudy (anonymous):

cos(2arccos1/3)=? please can you answear?

zepdrix (zepdrix):

\[\large \cos\left(2\color{royalblue}{\arccos\frac{1}{3}}\right)\] \[\large \color{royalblue}{\arccos\frac{1}{3}=\theta} \qquad\rightarrow\qquad \cos \theta=\frac{1}{3}=\frac{adjacent}{hypotenuse}\] We can draw a triangle to express this Trig function.

zepdrix (zepdrix):

Since this is your first question I feel I should ask, Is the formatted math text showing up ok on your screen? If you're using Internet Explorer it might just show up as a bunch of ugly code.

OpenStudy (anonymous):

no no thanks a lot bro

zepdrix (zepdrix):

hmm? :o

zepdrix (zepdrix):

Anyway, we can label the sides of our triangle given cosine of the angle.Then we want to use the Pythagorean Theorem to solve for the missing side.